Upon combustion, an unknown compound containing only carbon and hydrogen produces 40.29 g carbon dioxide and 16.49 g water. Determine the empirical formula of the unknown compound.
Answer #1

Now, on to answering the question. You need to find the MOLES of C, and H in order to find the empirical formula.

moles C = 40.29 g CO2 x 1 mole CO2/44 g x 1 mole C/mole CO2 = 0.92 moles C

moles H = 16.49 g H2O x 1 mole H2O/18 g x 2 = 1.83 moles H


C: 0.92 / 0.92 = 1.00
H: 1.83 / 0.92 = 2.00


You get a C/H mole ratio of 1/2. So the empirical formula is CH2.